Consider the titration of 25.00 mL of 0.082 M KI with 0.051 M AgNO3.

            Ksp for AgI is 8.3 X 10−17

Write the chemical reaction occurring in the beaker during the titration.

Calculate the equivalence point volume. (about 40 mL, calculate to 4 significant figures)

Calculate the pAg

a) Before AgNO3 is added (Boom, Why?)

            b) At the volume half way to the equivalence point (14.4)

            c) At the equivalence point (8.04)

            d) 5 mL after the equivalence point. (2.4)
Answer #1

The chemical reaction occurring in the beaker during the titration can be written as follows.

AgNO3 + KI AgI + KNO3

At equivalence point, (M*V)KI = (M*V)AgNO3

i.e. 0.082 M * 25 mL = 0.051 M * (V)AgNO3

Therefore, the equivalence point volume (V)AgNO3 = 40.20 mL